Answer the following problems


e. Consider a given bond that has five years maturity, Br.1000 face value and a 12 percent coupon rate. Suppose a broker’s commission of Br.50 is imposed by brokers to buy or sell the bond. Assume further, that the discount rate (minimum required rate of return) is 10 percent and the bond pays interest annually. What is the price of the bond?



f. Project X requires an immediate investment of $150,000 and will generate net cash inflows of $60,000 for the next three years. The project’s discount rate is 7%. If net present value is used to appraise the project, should Project X be undertaken?

1
Expert's answer
2022-03-08T08:51:57-0500

1.) Price of bond.


"Formula=C\times\frac{1-(1+r)^{-n}}{r}+\frac{FV}{(1+r)^n}"


Where; C is coupon payment="Coupon rate\times FV=.12\times1050=126"

Commission will increase face value.

Face Value=1,000+50=1,050


"Formula=126\times\frac{1-(1+0.1)^{-5}}{0.1}+\frac{1050}{(1+0.1)^5}=1,130"


Price of bond=Br. 1,130.


2.) Net Present Value(PV- Cost).


"Year 0=-150,000\times\frac{1}{(1.07)^{0}}=-150,000"


"Year 1=60,000\times\frac{1}{(1.07)^{1}}=56,075"


"Year 2=60,000\times\frac{1}{(1.07)^{2}}=52,406"


"Year 3=60,000\times\frac{1}{(1.07)^{3}}=48,978"


Sum all the above to find NPV.

NPV=Year 0+ Year 1+ Year 2+ Year 3

NPV=-150,000+56,075+52,406+48,978=$7,459